OpenStudy (anonymous):

Solve for X. 1/2 * X = 9

OpenStudy (anonymous):

Hey try multiply both sides by 2, what happens?

pooja195 (pooja195):

Multiply by the reciprocal 1/2 ---> 2/1 9 x 2=?

OpenStudy (michele_laino):

here I can multiply both sides by 2, so I get: \[\frac{2}{2} \cdot x = 2 \cdot 9\]

OpenStudy (anonymous):

18

pooja195 (pooja195):

Well done

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Yay

OpenStudy (anonymous):

:D

pooja195 (pooja195):

:D

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

so the answer is 18 yes?

pooja195 (pooja195):

Correct

OpenStudy (anonymous):

thx guys see u tamarrow

pooja195 (pooja195):

bai bai

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } x = 9 \implies \color{red}{2} \times \frac{ 1 }{ 2 }x = 9 \times \color{red}{2}\] notice when you multiply by 2 our expression here on the left side becomes 1 so we're left with just x since 1*x = x right? And then on the right we have 18 as what we do to one side, we must also do it to the other :)

OpenStudy (anonymous):

bai batman and pooja

OpenStudy (anonymous):

Take care! :)